### v1

# Learning in Structured Stackelberg Games

Abstract
We study structured Stackelberg games, in which both players (the leader and the follower) observe information about the state of the world at time of play. Importantly, this information may contain information about the follower, which the leader may use when deciding her strategy. Under this setting, we show that no-regret learning is possible if and only if the set of mappings from contexts to follower types that the leader uses to learn is not “too complex”. Specifically, we find that standard learning theoretic measures of complexity do not characterize learnability in our setting and we give a new dimension which does, which we term the Stackelberg-Littlestone dimension. In the distributional setting, we give analogous results by showing that standard complexity measures do not characterize the sample complexity of learning, but a new dimension called the Stackelberg-Natarajan dimension does. We then show that an appropriate empirical risk minimization procedure achieves the corresponding sample complexity.
1 Introduction
Many strategic interactions between a principal and an agent can be cast as Stackelberg games [36]. In a Stackelberg game, the principal (also called the leader) commits to a possibly randomized strategy, before the agent (also called the follower) observes this strategy and best-responds to it. Due to their effectiveness in modeling such leader-follower interactions, Stackelberg games have had significant real-world impact over the past decade, particularly in security domains [35]. For example, in protecting endangered species [38], a wildlife agency determines ranger patrol locations before poachers select targets. Coast guards decide on shoreline patrol routes in advance of illegal fishers’ actions [12] and, similarly, airport security [32] chooses checkpoint screening strategies before smugglers decide which entry points to use.
In many of these real-world settings, the state of the world at time of play can greatly influence the utilities of the players and subsequently the outcome of the game. For example, defending some areas of the airport might be deemed of higher- or lower-importance based on the schedule of flights and the number of passengers on board. Similarly, weather patterns might make it harder to reach animals in certain parts of the park, while breeding seasons might make others more vulnerable thereby affecting their risk levels as targets. When contextual information becomes available over time, using it could help the leader design more effective strategies against the follower. This raises a natural question: given that the state of the world might change in current and future rounds of play, how should the leader in a Stackelberg game design a policy that adapts to new contexts coming in over time?
One way is to study repeated Stackelberg games with side information [24], where at each interaction both the leader and the follower get to observe a context that affects their utilities. As in Balcan et al. [5], in this setting each follower that the leader faces might have different preferences and objectives, i.e. there might be different follower types. Harris et al. [24] study this problem and show that as long as either the followers or the contexts are independently drawn from a fixed and unknown distribution, no-Stackelberg-regret learning is possible. That is, the leader is guaranteed to achieve nearly as much utility as the best fixed policy in hindsight. In the fully adversarial regime, however, regret can grow linearly with the time horizon. Interestingly, this lower bound follows from the adversary’s ability to encode an arbitrarily complex relationship between the sequence of contexts and the follower types.
In this work, we take a more granular view of this impossibility result and ask whether introducing additional structure between the contextual information and the follower types can enhance the leader’s ability to maximize returns, even when both are chosen adversarially to the leader. Namely, there could be correlation between the two, making the context predictive of the type of follower about to appear next. In wildlife protection for example, specific tire tracks might signal the presence of a particular poaching group and market demand trends might be indicative of which animals’ poachers are now more likely to appear. Similarly, in airport security, current affairs often raise or lower the risk of an attack on certain international flights.
Algorithms proposed by Harris et al. [24] and later work [9] are designed to handle changes in utility caused by the context, but crucially they assume no structure in how contexts relate to follower types. If, however, there is a fixed and unknown mapping between contexts and follower types, can the leader learn to exploit it to achieve sublinear regret? In this case, the leader m ust adapt her policy based on which follower types are more likely to appear, given the context and past observations. This poses an additional and possibly more complex learning task for the leader, which we study in this work.
1.1 Our Contributions
In this paper, we revisit the framework of Harris et al. [24] and take a learning theoretic approach to analyzing the leader’s task in predicting the follower type from contexts. Importantly, we find that existing notions of complexity do not capture the difficulty of learning in this setting—we obtain (tight) sample complexity and Stackelberg regret guarantees that are parameterized by new notions of complexity and that directly exploit the structure of Stackelberg games. Concretely, our main contributions in this work are as follows:
  1. 1.
    We formalize online learning in Stackelberg games with side information, under the structural assumption that there exists a fixed, unknown mapping between contexts and follower types. In particular, we study how the leader, given access to a set of possible such mappings, can leverage it to learn an optimal policy. In particular, we study this problem under two different settings: (1) the realizable setting, in which there exists an optimal predictor of the followers’ types that the leader competes with; and (2) the agnostic setting, in which no such assumption is made.
  2. 2.
    We circumvent the impossibility result shown in Harris et al. [24] by providing algorithms with no-regret guarantees under the above structural assumption. We find that our problem is closely related to the problem of multiclass prediction, yet, interestingly, we show that our setting can be easier to learn in. In particular, we show that the (multiclass) Littlestone dimension, which quantifies the difficulty of learning in online multiclass prediction, can give loose mistake bounds in our setting. Motivated by this observation, we introduce a new notion of complexity, the Stackelberg-Littlestone dimension. We show that it characterizes learnability in our setting, and prove it to be at most as large as the multiclass Littlestone dimension. In fact, we give an example in which the latter is infinite, yet the former is zero. We also provide no-regret algorithms both for both the realizable and agnostic settings.
  3. 3.
    We introduce and formalize the study of distributional learning in Stackelberg games with side information. When the leader has access to a set of (context, follower type) examples, we show generalization with PAC-style bounds. Here too, we find that existing measures which characterize the complexity of multiclass prediction, such as the Natarajan dimension, do not characterize the sample complexity of our problem. Analogous to the online setting, we introduce the notion of Stackelberg-Natarajan dimension, with which we achieve an upper bound that is tight up to logarithmic factors.
1.2 Related Work
Stackelberg games. In the one-shot version of a Stackelberg game, Conitzer and Sandholm [16] prove that the problem of computing a Stackelberg equilibrium is NP-hard and they give an algorithm for maximizing the leader’s payoff. Other works focus on relaxing the assumption that all of the game’s parameters are known (e.g. utilities, number of actions) and aim to learn them instead [2, 31].
Our work is closely related to Bayesian Stackelberg games, in which the leader is playing against a single unknown follower type that is drawn from a given set of follower types [16]. This has been studied both in one-shot games, in which the leader must find a single optimal-mixed strategy [16, 30, 37] and in the repeated setting, in which the leader faces the same follower type sequentially and must learn from previous interactions [28, 26].
Learning in Stackelberg games. The problem of learning an optimal strategy for the leader with a single unknown follower type has been studied by Letchford et al. [26] and Blum et al. [11]. These results have been extended by Peng et al. [31], who give faster rates and almost-matching lower bounds. Recently, Bacchiocchi et al. [2] relax some structural assumptions from previous work, including tie-breaking rules for the follower, and revise worst-case sample complexity results of learning the leader’s optimal mixed strategy. Recent research has explored many domains through the framework of Stackelberg games, including strategic classification, where classifiers account for individuals who strategically modify their features [23]; Bayesian persuasion, which models how a leader strategically reveals information to influence a follower’s decisions [25]; interactions with calibrated followers, who update their strategies based on past observations rather than making immediate best responses [21]; and settings with non-myopic followers, who optimize their actions over multiple time steps rather than reacting myopically [20, 15].
Part of our work focuses on learning in online Stackelberg games. Online Stackelberg games with multiple follower types were introduced in Balcan et al. [5]. The authors show that when a series of unknown attacker types is chosen adversarially, playing the well-known Hedge algorithm  [19] over a special set of mixed strategies yields a no-regret algorithm.
Our departing point is the work of Harris et al. [24], in which the authors study the problem of online learning in Stackelberg games with side information. The authors show that when either the sequence of follower types or the sequence of contexts are drawn independently from some fixed, unknown distribution, no-regret learning is indeed possible. However, when both of them are chosen adversarially, regret can scale linearly in the worst case, as the problem can be reduced from online learning linear thresholds. On the other hand, we make a structural assumption about the relationship between contexts and follower types: namely that it belongs to a class of functions that the leader has access to. With this, we show that learning is possible, and our guarantees depend on an appropriate measure of complexity of this function class.
Multiclass learnability. Our work can also be viewed through the lens of multiclass learnability, as it suffices for the leader to learn the hypothesis class of the function mapping the contexts to a (finite) set of follower types. Then the leader can compute the optimal strategy against the predicted follower. Multiclass prediction has been studied extensively both in the ditributional ([29, 17, 13]) and in the online setting ([17, 22]). Indeed we adopt several tools from online multiclass prediction in our work. Interestingly, however, we show that learning in Stackelberg games with contexts is at most as hard as multiclass prediction, and in some cases, can be much easier. For example, in the online setting, we give an example in which the leader can learn to play optimally, under a hypothesis class that has infinite Littlestone dimension.
2 Problem Setting
Notation.
We use $[N]:=\{1,2,\ldots,N\}$ to denote the set of natural numbers up to and including $N$. We use $\Delta(\pazocal{A})$ to denote the probability simplex over a (finite) set $\pazocal{A}$.
Stackelberg games with side information.
Consider a game between a leader and a follower. Before the game starts, some contextual information $\mathbf{z}\in\pazocal{Z}\subseteq\mathbb{R}^{d}$ is revealed to both players. The leader plays first and commits to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$, where $\pazocal{A}$ is the set of finite actions available to the leader. The follower best responds to the leader’s strategy, $\mathbf{x}$, and the context, $\mathbf{z}$, by playing some action $a_{f}\in\pazocal{A}_{f}$, where $\pazocal{A}_{f}$ is the finite set of follower actions. [^2] Note that without loss of generality the follower’s best response will always be a pure strategy. We define the follower’s best response as:
$b_{f}(\mathbf{z},\mathbf{x})\in\arg\max_{a_{f}\in\pazocal{A}_{f}}\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u_{f}(\mathbf{z},a_{l},a_{f}).$ (1)
Here, $u_{f}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ is the utility function of follower $f$. We assume that the follower breaks ties in favor of the leader.
We allow the follower to be one of $K\in\mathbb{N}_{\geq 1}$ follower types in the set $\mathscr{K}=\{f^{(1)},\dots,f^{(K)}\}$. The type of follower characterizes their utility function, $u_{f^{(i)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$. We assume that the leader knows the set of possible follower types a priori, but not the realized type. The leader aims to maximize her own utility, $u:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$. By slight abuse of notation, we sometimes use the shorthand
$u(\mathbf{z},\mathbf{x},a_{f})=\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z},a_{l},a_{f}).$
We denote a Stackelberg game with side information as a tuple, $\pazocal{G}=(\pazocal{A},\pazocal{A}_{f},\pazocal{Z},u,u_{f^{(1)}},\ldots,u_{f^{(K)}})$, consisting of the leader’s action set, the follower’s action set, the set of possible contexts, as well as the leader’s and followers’ utility functions.
Remark 2.1.
For the remainder of the paper and without loss of generality, we will consider $\mathscr{K}$ to be a set of distinct follower types. We call any two follower types distinct, if the optimal leader’s strategy against them is not the same for every context.
3 Online Learning in Structured Stackelberg Games
Consider an online Stackelberg game in which the leader is faced by a sequence of followers. In each round $t\in[T]$, a context $\mathbf{z}_{t}\in\pazocal{Z}$ is revealed to both players. A follower of type $f_{t}$ arrives, the leader commits to a mixed strategy $\mathbf{x}_{t}\in\Delta(\pazocal{A})$, and the follower best responds to it and the context according to Equation 1. As before, we assume that the leader knows the utilities of all follower types, but the type of follower at time $t$ is not revealed to the leader until after the round is over. In this section, we make no assumptions about either the contexts or the follower types that appear at each timestep. In fact, we handle the case, where they both could be chosen jointly by an adversary.
3.1 Leader’s task in the realizable setting
Consider a set $\pazocal{H}$ (also called a hypothesis class) of candidate functions mapping contexts to follower types, which the leader uses to learn. More formally, for each $h\in\pazocal{H}$, we let $h:\pazocal{Z}\rightarrow\mathscr{K}$. We first consider the realizable setting, where there exists a true mapping $h^{*}\in\pazocal{H}$ such that $h^{*}(\mathbf{z}_{t})=f_{t}$ for all possible $(\mathbf{z}_{t},f_{t})$ pairs. We relax this assumption in Section 3.2. We use $h^{*}$ as a benchmark to measure the leader’s performance over time, and we quantify this via the notion of contextual Stackelberg regret.
Definition 3.1 (Contextual Stackelberg Regret).
Given a sequence of followers $f_{1},\ldots,f_{T}$ and contexts $\mathbf{z}_{1},\dots\mathbf{z}_{T}$, the leader’s contextual Stackelberg regret is given by:
$R(T):=\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})).$
We let $\pi_{h^{*}}(\mathbf{z})=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$ and we assume that there is a fixed and arbitrary tie-breaking rule over the leader strategies. We let $\mathbf{x}_{t}$ for $t\in[T]$ be the mixed strategy played by the leader at time $t$.
The goal of the leader is to learn a policy $\pi:\pazocal{Z}\to\Delta(\pazocal{A})$ that minimizes her regret, that is, it achieves utility comparable to the utility she would get if she could (1) optimally predict the follower type given the context and (2) play an optimal strategy against that follower type. In the remainder of this section, we will present an algorithm that meets that benchmark. In general, if the regret of an algorithm is sublinear in the time horizon $T$, i.e., $R(T)=o(T)$, we say that the algorithm is a no-regret algorithm.
Remark 3.2.
An important part of our analysis is that the leader be able to compute the optimal strategy for a fixed context and follower type. However, due to the tie-breaking rules of the follower, the leader might not be able to get an exact solution. Previous work [5, 24] shows that the leader can get an arbitrarily small error in the utility by optimizing over an appropriately selected subset of the simplex. For ease of notation, in the remainder of this paper, we will assume that the leader can directly optimize over the entire simplex, without that affecting her regret guarantees. For completeness, we formally state this result in Appendix A.1.
Background on Online Multiclass Learnability.
Our results build off of tools and techniques from multiclass online learning [17]. In this section, we highlight the concepts and definitions from this area that are relevant to our work. We start with the notion of shattering, which captures the following intuition: the more “complex” a set of functions is, the more complex patterns it should be able to fit.
Definition 3.3 ((Multiclass) $\pazocal{H}$-Shattered Littlestone Tree).
Let a Littlestone tree, $T$, be a tree whose internal nodes are instances in $\pazocal{Z}$ and where for each node the node-to-child edges correspond to different labels in a finite set $\mathscr{K}$. We say that $T$ is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a function $h\in\pazocal{H}$ such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $h(\mathbf{z}_{i})$.
The notion of multiclass Littlestone dimension reflects how the ability of a function class to shatter deeper trees corresponds to greater complexity in the class. [^3] Note that for binary labels, the definition of Littlestone dimension recovers the definition of the well-studied VC-dimension in the case where all nodes on the same depth of a shattered tree are the same. This definition is due to Daniely et al. [17], and corresponds to a natural generalization of the classic definition of the Littlestone dimension for binary classification [27] to the multiclass setting.
Definition 3.4 ((Multiclass) Littlestone Dimension).
The Littlestone dimension of a multiclass hypothesis class $\pazocal{H}$, denoted as $\textnormal{Ldim}(\pazocal{H})$, is the maximal integer $d$, such that there exists a full binary tree of depth $d$ that is shattered by $\pazocal{H}$.
Examples.
The finiteness of the Littlestone dimension of a hypothesis class is necessary for the class to be online learnable [27]. There are several examples of hypothesis classes, which have finite Littlestone dimension and can, therefore, be (online) learned. These include the class of conjunctions, the class of decision lists, and the class of linear thresholds with a margin (see Proposition 17 in [1]).
In Daniely et al. [17], the authors show that the multiclass Littlestone dimension of a hypothesis class characterizes the learnability of the class. In our setting, learning how to optimally predict the follower’s type given a context is an instance of multiclass prediction and is indeed sufficient for learning an optimal policy. Namely, given the context and the predicted follower type, the leader can deterministically find a strategy that maximizes her utility, by solving a sequence of $|\pazocal{A}_{f}|$ linear programs [16]. However, it is unclear a priori whether learning to optimally predict the follower’s type is also necessary. In Theorem 3.5 we show that it is not.
Theorem 3.5.
There exists a Stackelberg game, $\pazocal{G}$, and a hypothesis class, $\pazocal{H}$, such that $\textnormal{Ldim}(\pazocal{H})=\infty$, but the leader’s optimal policy can be determined without learning.
Proof.
Consider the setting where the context space is $\pazocal{Z}=\{1,2\}\times[0,1]$ and we have 4 follower types $\{f^{(1)},f^{(2)},f^{(3)},f^{(4)}\}$. Follower types $f^{(1)}$ and $f^{(2)}$ have fixed utility functions $u_{f^{(1)}},u_{f^{(2)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ such that for all $\mathbf{z}\in\pazocal{Z}$, we have that $\pi_{*}^{(1)}(\mathbf{z})\neq\pi_{*}^{(2)}(\mathbf{z})$, where $\pi_{*}^{(i)}$, refers to the optimal policy against follower $f^{(i)}$. Let follower type $f^{(3)}$ have utility function $u_{f^{(3)}}(\mathbf{z},a_{l},a_{f})=\mathbbm{1}\{\mathbf{z}[1]=1\}\cdot u_{f^{(1)}}(\mathbf{z},a_{l},a_{f})+\mathbbm{1}\{\mathbf{z}[1]=2\}\cdot u_{f^{(2)}}(\mathbf{z},a_{l},a_{f})$ and follower type $f^{(4)}$ have utility function $u_{f^{(4)}}(\mathbf{z},a_{l},a_{f})=\mathbbm{1}\{\mathbf{z}[1]=2\}\cdot u_{f^{(1)}}(\mathbf{z},a_{l},a_{f})+\mathbbm{1}\{\mathbf{z}[1]=1\}\cdot u_{f^{(2)}}(\mathbf{z},a_{l},a_{f})$. Let
$h_{\gamma_{1},\gamma_{2}}(\mathbf{z})=\begin{cases}f^{(1)}\;\;&\text{ if }0\leq\mathbf{z}[2]<\gamma_{1}\\ f^{(2)}\;\;&\text{ if }\gamma_{2}<\mathbf{z}[2]\leq 1\\ f^{(3)}\;\;&\text{ if }(\mathbf{z}[1]=1\text{ and }\gamma_{1}\leq\mathbf{z}[2]<\frac{1}{2})\text{ or }(\mathbf{z}[1]=2\text{ and }\gamma_{2}\leq\mathbf{z}[2]\leq 1/2)\\ f^{(4)}\;\;&\text{ if }(\mathbf{z}[1]=2\text{ and }\gamma_{1}\leq\mathbf{z}[2]<\frac{1}{2})\text{ or }(\mathbf{z}[1]=1\text{ and }\gamma_{2}\leq\mathbf{z}[2]\leq 1/2)\\ \end{cases}$
[…]
Figure 1: Follower type as a function of the context $\mathbf{z}$ in the example of Theorem 3.5. Observe that while $\gamma_{1}$ and $\gamma_{2}$ vary with each hypothesis, the optimal mixed strategy to play is determined by whether $\mathbf{z}[2]\leq\frac{1}{2}$.
Theorem 3.5 shows that the Littlestone dimension does not characterize the complexity of online learning in Stackelberg games with side information. The example constructed in our proof highlights that the Littlestone dimension is blind to how the hypothesis class interacts with the utilities of the players. This observation motivates the definition of a new quantity, which we term the Stakelberg-Littlestone (SL) dimension. In order to define the SL dimension, we first need to introduce the notion of an SL-Tree, which is a natural extension of the Littlestone tree to our setting.
Consider a tree of depth $d$ whose internal nodes are instances in $\pazocal{Z}$ and where the node-to-child edges for a single node, correspond to different labels in a finite set $\mathscr{K}$. The tree $T$ is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a function $h\in\pazocal{H}$ such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $h(\mathbf{z}_{i})$.
We say that $T$ is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a function $h\in\pazocal{H}$ such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $h(\mathbf{z}_{i})$.
Definition 3.6 ($\pazocal{H}$-Shattered Stackelberg-Littlestone Tree).
For a Stackelberg game, $\pazocal{G}$, let a Stackelberg-Littlestone tree, $T$, be a tree whose internal nodes are instances in $\pazocal{Z}$ and where for each node, the node-to-child edges correspond to different vectors in $\Delta(\pazocal{A})$. We say that $T$ is shattered by a hypothesis class $\pazocal{H}$ with respect to $\pazocal{G}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\ldots,\mathbf{z}_{n}$, there exists a hypothesis $h\in\mathscr{H}$, such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $\mathbf{x}_{i}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{i},\mathbf{x},b_{h(\mathbf{z}_{i})}(\mathbf{z}_{i},\mathbf{x}))$. In the case of ties, we assume that the leader follows a fixed and arbitrary tie-breaking rule.
Equipped with these definitions, we are now ready to introduce the Stackelberg-Littlestone dimension, which characterizes the complexity of online learning in our setting.
Definition 3.7 (Stackelberg-Littlsestone Dimension).
The Stackelberg-Littlestone dimension of a hypothesis class $\mathscr{H}$ with respect to a Stackelberg game, $\pazocal{G}$, denoted as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, is the maximal integer $d$, such that there exists a full binary SL tree of depth $d$ that is shattered by $\pazocal{H}$.
3.1.1 Lower bound
We begin by showing that for any online Stackelberg game $\mathscr{G}$, no algorithm can achieve regret smaller than $\Omega(\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H}))$.
Theorem 3.8.
There exists a Stackelberg game, $\pazocal{G}$, such that for any hypothesis class $\pazocal{H}$ any deterministic algorithm will suffer at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ regret. Furthermore, any randomized algorithm will suffer at least $\frac{1}{2}\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ regret in expectation.
Proof.
Consider a Stackelberg game, in which for any fixed context, playing the optimal strategy against the true follower type yields utility $1$, whereas playing any other strategy yields utility 00. Concretely, let $\beta_{1},\beta_{2},\ldots,\beta_{K}$ be distinct real numbers and consider the following payoff matrix, where the leader is the row player and the column player is follower type $f^{(i)}$. The leader has 2 pure actions and the follower has 3 pure actions.
$F_{1}$ $F_{2}$ $F_{3}$
$L_{1}$ $0,\frac{1}{\beta_{i}}$ $1,1$ $0,0$
$L_{2}$ $0,0$ $1,1$ $0,\frac{1}{1-\beta_{i}}$
Notice that if the leader puts probability $p>\beta_{i}$ on action $L_{1}$, then the follower will play action $F_{1}$ and the leader gets payoff 00. Similarly, if the leader puts probability $p<\beta_{i}$ on action $L_{1}$ then the follower plays action $F_{3}$ and the leader gets payoff 00. However, if the leader puts probability $p=\beta_{i}$ on action $L_{1}$, then the follower is indifferent and since the follower breaks ties in favor of the leader, she plays action $F_{2}$, giving the leader a payoff of $1$.
(Claim 1.) An adversary can force the learner to make $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes by using the following strategy: (1) select $\mathbf{z}_{t}$ as the next input from an SL-Tree of depth $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ that is shattered by $\pazocal{H}$ and (2) set the optimal mixed strategy, $\mathbf{x}_{t}^{*}$ to be the edge in the shattered tree not chosen by the leader, that is $\mathbf{x}_{t}^{*}\neq\widehat{\mathbf{x}}_{t}$, where $\widehat{\mathbf{x}}_{t}$ is the strategy played by the learner at time $t$. Since the tree is shattered by $\pazocal{H}$, there must exist some $h\in\pazocal{H}$ that is consistent with $(\mathbf{z}_{t},f_{t})_{t=1}^{\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})}$. Therefore, any deterministic learner will make at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
(Claim 2.) When the learner is randomized, at each timestep $t$, the adversary can select an $\mathbf{x}_{t}^{*}$ such that it minimizes the probability that the leader chooses $\widehat{\mathbf{x}}_{t}=\mathbf{x}_{t}^{*}$. In this case, the probability that the learner will make a mistake is at least 0.5, and equality holds when the leader selects uniformly among two branches of the tree. Summing over each layer of the tree gives us the lower bound on the learner’s expected number of mistakes. ∎
3.1.2 Upper bounds
Having established that the Stackelberg-Littlestone dimension characterizes the difficulty of learning in our setting, in this section we present an algorithm and show that it achieves optimal regret. Our guarantees in this setting are with respect to the leader’s number of mistakes (i.e., timesteps where the utility of the leader is not optimal). By the realizability assumption and the boundedness of the utility function, all mistake bounds imply the same regret bounds.
Warm-up: context-free utilities.
As a warm-up to our main results, consider a simpler setting, in which both the leader and the follower utility functions do not take the context as an input (i.e. they are context-free). This means that the context at each timestep is now only predictive of the follower’s type at that timestep. In Theorem 3.9, we show that we can reduce the problem of learning the optimal policy of the leader to online multiclass prediction. Additionally, the known optimal algorithm in online multiclass prediction (the Standard Optimal Algorithm (SOA)) meets the lower bound shown in Theorem 3.8.
Theorem 3.9.
Consider a Stackelberg game, $\pazocal{G}$, such that the utilities of both the leader and the follower are context-free. Then for any finite class, $\pazocal{H}$, of hypotheses $h:\pazocal{Z}\to\mathscr{K}$, there exists a reduction of this setting to multiclass prediction that runs in $O(\text{poly}(K,|\pazocal{A}_{f}|,|\pazocal{A}|,|\pazocal{H}|))$ time and achieves regret at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
Proof.
The reduction is as follows. First, compute the optimal strategy against each type of follower offline. This can be achieved using linear programming and requires $O(\text{poly}(K,|\pazocal{A}_{f}|,|\pazocal{A}|))$ runtime [16]. Then create a (possibly many-to-one) look-up table mapping each follower type $f^{(i)}$ to an optimal strategy, $\mathbf{x}_{*}^{(i)}$, for the leader against that follower. Compute $\pazocal{H}^{\prime}=\{h^{\prime}(\mathbf{z}):=\sum_{i=1}^{K}\mathbf{x}_{*}^{(i)}\cdot\mathbbm{1}\{h(\mathbf{z})=f^{(i)}\}\}_{h\in\pazocal{H}}$, which takes at most $O(\text{poly}(|\pazocal{H}|))$ time. Finally, we run the multiclass Standard Optimal Algorithm (see [17]) on $\pazocal{H}^{\prime}$ considering every optimal strategy for the leader in the look-up table as one label. It remains to show that $\textnormal{Ldim}(\pazocal{H}^{\prime})=\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$. Notice that $\pazocal{H}^{\prime}$ is the set of all hypotheses, mapping each context to an optimal strategy for the leader against every follower type that is consistent with $\pazocal{H}$ for that context. Consider any SL tree and any Littlestone tree such that the context at each corresponding node is the same among them. If the SL tree is shattered by $\pazocal{H}$ then for every root-to-leaf path there must exist a hypothesis in $\pazocal{H}$ that is consistent with the labels along that path. But by definition of $\pazocal{H}^{\prime}$ for that corresponding path in the Littlestone tree there must exist an $h^{\prime}\in\pazocal{H}^{\prime}$ that is consistent with the labelings of that path. With the same argument, the converse is also true. Therefore, the SL tree with maximal depth that is shattered by $\pazocal{H}$ is precisely the generalized Littlestone tree with maximal depth that is shattered by $\pazocal{H}^{\prime}$. ∎
We note that the algorithm described above coincides with Algorithm 1 (presented below), which we prove to be optimal in the general contextual Stackelberg game setting.
Context-dependent utilities.
We now let the utilities of both players to be functions of the context. We present our main algorithm (Algorithm 1), which we show achieves regret that matches the lower bound in Theorem 3.8 for any Stackelberg game. Our approach is inspired by the Standard Optimal Algorithm for online multiclass prediction, with the main difference that our algorithm uses a hypothesis class, $\pazocal{H}$, to optimize directly over leader strategies in the simplex.
Intuitively, our algorithm keeps track of all the “surviving” hypotheses in $\pazocal{H}$, that is, the hypotheses that are consistent with the data seen so far. At each timestep the algorithm observes the context and considers the predicted follower types of all surviving hypotheses. For every such follower type the algorithm computes the optimal leader strategy against them. Then it plays the strategy that maximizes the Stackelberg Littlestone dimension of the set of surviving hypotheses were that to be the optimal strategy for that timestep. If the algorithm makes a mistake, then the set of surviving hypotheses will have strictly smaller Stackelberg-Littlestone dimension than before. In other words, the algorithm learns from this mistake. Notice that the number of possible strategies in the simplex is infinite, yet our algorithm only considers a bounded number of them at each timestep, at most one for each follower type.
Algorithm 1 Stackelberg Standard Optimal Algorithm (SSOA)
1:input: time horizon, $T$, Stackelberg game $\pazocal{G}$, and hypothesis class $\pazocal{H}$, where $h:\pazocal{Z}\to\mathscr{K}$ for every $h\in\pazocal{H}$.
2:Set $V_{0}=\pazocal{H}$
3:for $t\in[T]$ do
4:     Observe $\mathbf{z}_{t}$
5:     Let $\mathbf{x}^{(h,t)}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))\quad\forall h\in\pazocal{H}$
6:     Let $X_{t}=\{\mathbf{x}^{(h)}:h\in V_{t-1}\}$triangleright Note that $|X_{t}|\leq K$
7:     Let $V_{t}^{(\mathbf{x})}=\{h\in V_{t-1}:\mathbf{x}^{(h,t)}=\mathbf{x}\}\quad\forall\mathbf{x}\in X_{t}$
8:     Play action $\mathbf{x}_{t}=\arg\max_{\mathbf{x}\in X_{t}}\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$
9:     Observe $f_{t}$
10:     Set $V_{t}=\{h\in V_{t-1}:h(\mathbf{z}_{t})=f_{t}\}$
11:end
Theorem 3.10.
For any Stackelberg game, $\pazocal{G}$ and every hypothesis class, $\pazocal{H}$, the regret of Algorithm 1 is at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
Proof.
Let $T_{e}\subseteq[T]$ be the set of time steps on which SSOA errs and let $T_{c}=[T]\setminus T_{e}$ be the set of time steps on which SSOA correctly predicts the optimal strategy against the true follower type. Recall from Remark 3.2 that we can make $\sum_{t\in T_{c}}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$ arbitrarily small.
Now from the boundedness of the utility function, we also know that $\sum_{t\in T_{e}}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))\leq|T_{e}|$. It suffices to show that for every time step $t\in T_{e}$ that SSOA makes a mistake, we have that $\textnormal{SLdim}_{\pazocal{G}}(V_{t})\leq\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})-1$. For the sake of contradiction, assume that $\textnormal{SLdim}_{\pazocal{G}}(V_{t})=\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})$. Given that SSOA chooses a strategy $\mathbf{x}$ that maximizes $\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$ and that SSOA made a mistake at this timestep, there must have been at least two distinct strategies $\mathbf{x},\mathbf{x}^{\prime}$ such that $\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})=\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x}^{\prime})})=\textnormal{SLdim}_{\pazocal{G}}(V_{t})$. However, this would mean that we could construct a full binary SL tree that is shattered by $\pazocal{H}$ and that has depth $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})+1$, which leads to a contradiction.
3.1.3 Relationship with Littlestone dimension
We have so far shown that the Stackelberg-Littlestone dimension characterizes learnability in our setting. We next explore how that relates to multiclass learnability. While, as discussed, learning to predict the follower type from the context is a sufficient condition for learning the leader’s optimal policy, in this section we prove that this is not a necessary condition. In Theorem 3.11, we show that the Littlestone dimension is always at most as large as the Stackelberg Littlestone dimension and in Proposition LABEL:prop:ldim=sldim we show that equality always holds in the special case of two follower types. In Theorem 3.12, we show that an out-of-the-box application of the optimal algorithm for multiclass learning in our setting can result not only in suboptimal number of mistakes, but also in suboptimal utility for the leader. This is in contrast to our proposed algorithm, which achieves optimal utility by accounting for the interaction between the hypothesis class and the Stackelberg game.
Theorem 3.11.
For any hypothesis class $\pazocal{H}$, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$.
Proof.
If the leader plays SOA, she will make at most $\textnormal{Ldim}(\pazocal{H})$ mistakes. This is because she will predict the follower’s type incorrectly at most $\textnormal{Ldim}(\pazocal{H})$ times, and predicting the follower’s type correctly is a sufficient condition for playing the optimal mixed strategy. Since the SOA makes $\textnormal{Ldim}(\pazocal{H})$ mistakes and we know that any deterministic algorithm will make at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes by Theorem 3.8, we can infer that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$. ∎
While it is known that SOA makes at least $\textnormal{Ldim}(\pazocal{H})$ mistakes when learning a hypothesis class, $\pazocal{H}$, a careful reader might notice that this does not immediately show that SOA achieves suboptimal regret in our setting. This is because mispredicting the follower type does not necessarily imply that the optimal strategy against that and the true follower type do not yield the same utility for the leader. In Theorem 3.12, we prove that, indeed, there are instances where SOA will achieve smaller cumulative utility for the leader compared to our proposed algorithm.
Theorem 3.12.
There exists some hypothesis class $\pazocal{H}$ and a Stackelberg game, $\pazocal{G}$, such that running multiclass SOA on $\pazocal{H}$ yields suboptimal utility for the leader.
Proof.
Consider $\pazocal{Z}=\{1,2,3,4\}$ and $\mathscr{K}=\{f^{(1)},f^{(2)},f^{(3)}\}$. We define $\pazocal{H}=\{h_{\sigma}(z):=\mathbbm{1}\{z\neq 4\}\cdot f^{(\sigma(z))}+\mathbbm{1}\{z=4\}\cdot f^{(3)}\}_{\sigma\in\Sigma}\cup\{h,h^{\prime}\}$ , where $\Sigma$ is the set of all permutations of the set $\{1,2,3\}$ and $h,h^{\prime}$ are defined as follows.
$h(z)=\begin{cases}f^{(3)}\;\;&\text{ if }z=1\\ f^{(2)}\;\;&\text{ if }z=2\\ f^{(1)}\;\;&\text{ if }z=3\\ f^{(2)}\;\;&\text{ if }z=4\\ \end{cases}\hskip 56.9055pth^{\prime}(z)=\begin{cases}f^{(3)}\;\;&\text{ if }z=1\\ f^{(1)}\;\;&\text{ if }z=2\\ f^{(2)}\;\;&\text{ if }z=3\\ f^{(2)}\;\;&\text{ if }z=4\\ \end{cases}$
Notice that the Littlestone dimension of this hypothesis class is at least $2$. This follows from the $\pazocal{H}$-shattered tree construction shown in Figure 2.
Figure 2: Construction of an $\pazocal{H}$-shattered Littlestone tree. Each node represents a context in $\pazocal{Z}$ and each edge corresponds to a labeling of the parent node to a follower type.
Consider now a Stackelberg game with utility functions $u^{(1)},u^{(2)},u^{(3)}$ corresponding to follower types $1,2$, and $3$ respectively, defined such that if $z=1$ or $z=2$ then $\pi_{*}^{(1)}(z)=\pi_{*}^{(2)}(z)\neq\pi_{*}^{(3)}(z)$, if $z=3$ then $\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$, and if $z=4$, then $\pi_{*}^{(1)}(4)\neq\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$. Here, we let $\pi_{*}^{(i)}:\pazocal{Z}\to\Delta(\pazocal{A})$ denote the optimal leader’s policy against follower type $f^{(i)}$. Concretely, consider a Stackelberg game with the leader being the row player and each follower being the column player, as shown in Table 1.
$F_{1}$ $F_{2}$
$L_{1}$ $2,1$ $0,0$
$L_{2}$ $0,0$ $2,0$
(a) Follower type $f^{(1)}$
$F_{1}$ $F_{2}$
$L_{1}$ $2,\mathbbm{1}_{\{z\neq 4\}}$ $0,0$
$L_{2}$ $0,0$ $2,\mathbbm{1}_{\{z=4\}}$
(b) Follower type $f^{(2)}$
$F_{1}$ $F_{2}$
$L_{1}$ $2,\mathbbm{1}_{\{z=3\}}$ $0,0$
$L_{2}$ $0,0$ $2,\mathbbm{1}_{\{z\neq 3\}}$
(c) Follower type $f^{(3)}$
Table 1: Utility tables for the leader (row player) and each follower type (column player)
We now show that the $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=1$. We will consider any possible SL tree. First, note that the context $z=4$ cannot be a node in an $\pazocal{H}$-shattered SL-Tree since for $z=4,\;\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$, where $f^{(2)}$ and $f^{(3)}$ are the only possible follower types predicted by hypotheses in $\pazocal{H}$ for this context. Similarly, $z=3$ cannot be a node in an $\pazocal{H}$-shattered SL tree, since the leader’s optimal strategy against all follower types is identical ($\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$). So it suffices to consider contexts $1$ and $2$. In Figure 3 we show the two $\pazocal{H}$-shattered SL tree constructions with $z=1$ and $z=2$ as the root nodes respectively. Notice that the maximal depth of a full binary tree is $1$, as desired.
Say that the adversary presents context $1$ followed by context $2$. We go through the SOA outputs per timestep below.
  • ($t=1$) This node has three children corresponding to follower types $f^{(1)},f^{(2)},f^{(3)}$, with $f^{(1)}$ and $f^{(2)}$ leading to a full binary subtree of maximal depth 1 each. To see this, notice that (1) neither $h$ nor $h^{\prime}$ are consistent with the mapping of context $z=1$ to follower type $f^{(1)}$ or $f^{(2)}$, and (2) $\pazocal{H}\setminus\{h,h^{\prime}\}$ consists of all permutations mapping contexts $\{1,2,3\}$ to follower types $\{f^{(1)},f^{(2)},f^{(3)}\}$. Therefore, at most two of these contexts can exist in any root-to-leaf path of any $\pazocal{H}$-shattered tree. The child corresponding to follower type $f^{(3)}$ leads to a full binary subtree of maximal depth at least $2$, as shown in Figure 2. So, the predicted follower type is $f_{1}=f^{(3)}$ and the strategy played is $x_{1}=\pi_{*}^{(3)}(1)$. However, the true follower type is $f_{1}^{*}=f^{(1)}$ (mistake).
  • ($t=2$) This node has two children mapping $z=2$ to follower types $f^{(2)}$ and $f^{(3)}$. No subsequent context can lead to a full-binary tree that is consistent with any of these assignments, so SOA tie-breaks. The adversary in this case can always force a mistake by claiming the correct follower type to be the one not predicted by the algorithm. So if SOA predicts $f_{2}=f^{(2)}$, and plays strategy $x_{2}=\pi_{*}^{(2)}(2)$ the adversary can choose the true follower type to be $f_{2}^{*}=f^{(3)}$ (mistake).
This leads to SOA mispredicting the follower type and getting suboptimal utility in both rounds. Since the Stackelberg Littlestone dimension is $1$ and we have shown that for the above sequence of contexts SOA makes $2$ mistakes, this concludes our proof.
For ease of exposition, we also provide an overview of SSOA for the same series of nodes (see Figure (3(a)) for reference).
  • ($t=1$) This root node has two branches: $\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and $\pi_{*}^{(3)}(1)$. Given the arguments used to prove that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=1$, the former branch leads to a full binary subtree of depth $1$, while the second branch cannot lead to a full binary tree. So SSOA chooses $x_{1}=\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and the true follower type is $f_{1}^{*}=f^{(1)}$ (no mistake).
  • ($t=2$) The context $z=2$ leads to two branches: $\pi_{*}^{(2)}(2)$ and $\pi_{*}^{(3)}(2)$, none of which leads to a complete binary tree. SSOA tie breaks, in which case an adversary can always force a mistake by choosing the correct follower type to be the one not predicted by SSOA (mistake).
Figure 3: $\pazocal{H}$-shattered SL-Tree with (a) $z=1$ as the root node and (b) $z=2$ as the root node.
While we have shown that running multiclass SOA alone can in certain settings achieve suboptimal utility for the leader, in the following theorem we show that our algorithm may be interpreted as running SOA in a transformed space.
Theorem 3.13.
For every online contextual Stackelberg game with $K$ distinct follower types and every hypothesis class $\pazocal{H}$, there exists a hypothesis class $\Pi$ such that running multiclass SOA on $\Pi$ incurs $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes. Moreover, this algorithm coincides with SSOA.
Proof.
Consider a mapping $\zeta:\pazocal{H}\to\Pi$ that maps every $h\in\pazocal{H}$ to a function $\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})$ such that:
$\pi_{h}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x})).$
Notice that $\zeta$ is a bijection given that each follower type is distinct (see Remark 2.1). Also note that given that learning $\pazocal{H}$ is a realizable problem, so is learning $\Pi$. This is because if there exists $h^{*}\in\pazocal{H}$ such that $\forall t\in[T]\quad h^{*}(\mathbf{z}_{t})=f_{t}$, then by definition there must exist $\pi^{*}\in\Pi$ such that $\forall t\in[T]\quad\pi^{*}(\mathbf{z}_{t})=\mathbf{x}^{*}_{t}$, since $\mathbf{x}^{*}_{t}=\arg\max_{\mathbf{x}\in\Delta(A)}u(\mathbf{z}_{t},\mathbf{x},b_{h^{*}(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))$.
We know that multiclass SOA makes at most $\textnormal{Ldim}(\Pi)$ mistakes for any hypothesis class $\Pi$ with infinite labels (see Theorem 1 in [22]). Therefore, it remains to show that $\textnormal{Ldim}(\Pi)=\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
Consider any SL-Tree for $\pazocal{H}$. Notice that all nodes are contexts and all edges represent optimal leader strategies, so this is a valid Littlestone tree for $\Pi$ as well. Also, for any root-to-leaf path on this tree that is consistent with a function $h\in\pazocal{H}$, there exists $\pi_{h}\in\Pi$ that is consistent with it as well, by definition of $\pi_{h}$. The converse is also true: in any Littlestone tree for $\Pi$ consider any root-to-leaf path, that is consistent with some $\pi_{h}\in\Pi$. Then by $\zeta$ being a bijection, there exists an $h\in\pazocal{H}$ that is also consistent with it. Therefore, a tree is SL-shattered by $\pazocal{H}$ if and only if it is shattered by $\Pi$. This holds for all trees, so it will hold for the maximal-depth trees as well.
3.2 No-regret learning in the agnostic setting
In this section we relax the realizability assumption, and we consider the setting where the sequence of follower types $f_{1},\cdots,f_{T}$ is not necessarily consistent with some hypothesis $h\in\mathscr{H}$. Unlike the realizable setting, here we will not have an absolute benchmark such as the mistake bound. Instead, we measure the performance of our algorithms relative to the best in hindsight hypothesis in the class. We define $h^{*}\in\arg\max_{h\in\mathscr{H}}\sum_{t\in[T]}\mathbbm{1}\{h(\mathbf{z}_{t})=y_{t}\}$ and generalize the notion of regret:
Definition 3.14 ((Agnostic Setting) Contextual Stackelberg Regret).
Given a sequence of followers $f_{1},\dots,f_{T}$ and contexts $\mathbf{z}_{1},\dots\mathbf{z}_{T}$, the leader’s contextual Stackelberg regret in the agnostic setting is given by:
$R(T):=\max_{h\in\mathscr{H}}\sum_{t\in[T]}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})),$
where $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$ is the sequence of mixed strategies played and $\pi_{h}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))$.
Combining our SSOA algorithm with results and tools from multiclass prediction [17], we get the following regret bound.
Theorem 3.15.
Consider a Stackelberg game, $\pazocal{G}$, and a hypothesis class, $\pazocal{H}$. In the agnostic setting, there exists an algorithm that achieves regret $R(T)\leq\sqrt{\frac{1}{2}\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})T\log{(TK)}}$.
We use a similar approach to Lemma 14 from [10] to get a (nearly-)matching lower bound.
Theorem 3.16.
Let $\pazocal{H}$ be any hypothesis class. There exists a Stackelberg game, $\pazocal{G}$, such that $\pazocal{H}$ has a finite Stackelberg-Littlestone dimension $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ and such that no algorithm in the agnostic, full-feedback setting can achieve regret smaller than $\Omega\left(\sqrt{\frac{1}{8}\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})T}\right)$.
We include the complete proofs of Theorems 3.15 and 3.16 in Appendix A.3.
4 Distributional Learning
We now shift our focus from online learning to distributional learning. Suppose that there exists an unknown, joint distribution over contexts and follower types. We first study the simpler setting, without contexts, in which we give strong generalization guarantees by exploiting the structure of the leader’s payoff function. When contexts come into play in Section 4.2, we analyze an intuitive approach that uses a black box empirical risk minimizer (ERM) and show that it is suboptimal in our setting. Analogous to our online results, we then give a better sample complexity bound that depends on a new complexity measure.
4.1 Learning without Contexts
Suppose there is a fixed, unknown distribution, $\mathscr{D}$, over the set of possible follower types. We denote with $u(\mathbf{x},b_{f^{(i)}}(\mathbf{x}))$ the leader’s payoff when committing to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$ against a follower of type $f^{(i)}$. We use the notation $u_{\mathscr{D}}(\mathbf{x})=\mathbb{E}_{f\sim\mathscr{D}}[u(\mathbf{x},b_{f}(\mathbf{x}))]$ and given a sample $S$ of follower types, we write:
$u_{S}(\mathbf{x})=\frac{1}{|S|}\sum_{f\in S}u(\mathbf{x},b_{f}(\mathbf{x}))$
Given a set $S$ of $m$ samples from $\mathscr{D}$, the goal of the leader is to choose a mixed strategy that fares well in expectation against any new sample from $\mathscr{D}$. In other words, we want to choose $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ such that with high probability:
$\mathbb{P}_{S}[\max_{\mathbf{x}}u_{\mathscr{D}}(\mathbf{x})-u_{\mathscr{D}}(\hat{\mathbf{x}})>\epsilon]\leq\delta.$
For this, it suffices to show a generalization guarantee, i.e. a bound on the difference between the leader’s expected payoff and the empirical payoff over a set of follower types for any possible mixed strategy.
Definition 4.1.
A generalization guarantee for the leader in a Stackelberg game, $\pazocal{G}$, is a function $\epsilon_{\pazocal{G}}:\mathbb{Z}_{\geq 1}\times(0,1)\to\mathbb{R}_{\geq 0}$ such that for any $\delta\in(0,1)$, any $m\in\mathbb{Z}_{\geq 1}$, and any distribution $\mathscr{D}$ over the follower types, with probability at least $1-\delta$ over the draw of a set $S\sim\mathscr{D}^{m}$ for any $\mathbf{x}\in\Delta(\pazocal{A})$, the difference between the average payoff of $\mathbf{x}$ over $S$ and the expected payoff of $\mathbf{x}$ over $\mathscr{D}$ is at most $\epsilon_{\pazocal{G}}(m,\delta)$:
$\mathbb{P}_{S\sim\mathscr{D}^{m}}[\exists\mathbf{x}\in\Delta(\pazocal{A})\;\;\text{such that}\;\;|u_{\mathscr{D}}(\mathbf{x})-u_{S}(\mathbf{x})|>\epsilon_{\pazocal{G}}(m,\delta)]\leq\delta$
The implication of a generalization guarantee is that it relates the utility of the leader’s strategy that maximizes the average payoff over the training data to the expected utility of the optimal strategy. Formally, this means that for $\hat{x}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{S}(\mathbf{x})$ and $x^{*}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{\mathscr{D}}(\mathbf{x})$:
$\mathbb{P}_{S\sim\mathscr{D}^{m}}[|u_{\mathscr{D}}(\mathbf{x}^{*})-u_{S}(\hat{\mathbf{x}})|>2\epsilon_{\pazocal{G}}(m,\delta)]\leq\delta.$
Consider the class $\pazocal{P}=\{u_{\mathbf{x}}:\mathscr{K}\to\mathbb{R}\;\;\text{such that}\;\;u_{\mathbf{x}}(f)=u(\mathbf{x},b_{f}(\mathbf{x}))\}$ to be the set of payoff functions for the leader. Each payoff function is parameterized by the leader’s mixed strategy, takes as input the follower’s type, and outputs the leader’s utility. We also define the dual class $\tilde{\pazocal{P}}=\{\tilde{u}_{f}:\Delta(\pazocal{A})\to\mathbb{R}\;\;\text{such that}\;\;\tilde{u}_{f}(\mathbf{x})=u(\mathbf{x},b_{f}(\mathbf{x}))\}$ to be the set of payoff functions for the leader, parameterized by a follower type that take as input the leader’s mixed strategy and output the leader’s utility. Consider the following definition, introduced by [6] and adapted to our setting:
Definition 4.2.
A class $\pazocal{H}$ of hypotheses, $h:\mathscr{K}\to\mathbb{R}$, is $(d,t)$-delineable if:
  1. 1.
    The class $\pazocal{H}$ consists of hypotheses that are parameterized by vectors $\mathbf{x}$ from a set $\pazocal{X}\subseteq\mathbb{R}^{d}$; and
  2. 2.
    For any $f\in\mathscr{K}$, there is a set $\mathscr{J}$ of $t$ hyperplanes such that for any connected component $\pazocal{X}^{\prime}=\pazocal{X}\setminus\mathscr{J}$, the dual function $h_{f}(\mathbf{x})$ is linear over $\pazocal{X}^{\prime}$.
With this definition, we make the following useful proposition.
Proposition 4.3.
The class of leader payoff functions, $\pazocal{P}$ that are parameterized by the leader’s mixed strategy and take as input the follower’s type is $(|\pazocal{A}|,\binom{|\pazocal{A}_{f}|}{2})$-delineable.
Proof.
The statement holds from the following two facts: (1.) Given a fixed follower type, $f$, we have that $\tilde{u}_{f}(\mathbf{x})$ is a piece-wise linear function. (2.) For any fixed follower type, $f$, we have that $u_{f}(\mathbf{x})$ will have at most $|A_{f}|(|A_{f}|-1)$ hyperplanes going through it. Each hyperplane corresponds to a change in best-response strategies for the follower from one pure action to another, so there can be at most $\binom{|\pazocal{A}_{f}|}{2}$ hyperplanes. ∎
Next, we present the definition of pseudo-dimension [33], which is complexity measure that characterizes a hypothesis class and in our setting helps us reason about how difficult the learning task of the leader is. Informally speaking, the more complex the hypothesis class, $\pazocal{P}$, the harder it is for the leader to distinguish between strategies, $\mathbf{x}$, which parameterize the class, and therefore, more samples are needed to learn.
Definition 4.4 (Pseudo-dimension).
Let $\pazocal{F}$ be a class of functions $f:A\to\mathbb{R}$, with an abstract domain, $A$. We say that $y^{(1)},\ldots,y^{(n)}$, witness the shattering of $S=\{x^{(1)},\ldots,x^{(n)}\}\subseteq A$ if for all $T\subseteq S$, there exists a function $f_{T}\in\pazocal{F}$, such that for all $x^{(i)}\in T,\;f_{T}(x^{(i)})\leq y^{(i)}$ and for all $x^{(i)}\not\in T,\;f_{T}(x^{(i)})>y^{(i)}$. We then say that the pseudo-dimension of $\pazocal{F}$, denoted $\textnormal{Pdim}(\pazocal{F})$, is the size of the largest set that can be shattered by $\pazocal{F}$.
Leveraging results from the data-driven algorithm design literature ([4, 8, 7]), such as Lemma 3.10 in [6] and Proposition 4.3 we have the following result:
Theorem 4.5.
The pseudo-dimension of the hypothesis class, $\pazocal{P}$ satisfies $\textnormal{Pdim}(\pazocal{P})\leq 9|\pazocal{A}|\log\left(4|\pazocal{A}||\pazocal{A}_{f}|^{2}\right)$.
So we also get the generalization guarantee from Pollard [33] and Dudley [18] with function:
$\epsilon_{\pazocal{G}}(m,\delta)=O\left(U\sqrt{\frac{|\pazocal{A}|\log(|\pazocal{A}||\pazocal{A}_{f}|^{2})}{m}}+U\sqrt{\frac{\ln(1/\delta)}{m}}\right),$
where $U$ is the maximum payoff value for the leader.
Discussion.
We note that the above generalization result is particularly strong. Informally, it asserts that no matter what the distribution over the follower types is, if the leader has so far encountered a number of followers that scales linearly with the size of her action set and logarithmically with the follower’s action set, it suffices for her to play a strategy that maximizes her average utility over the followers seen so far. This result holds even for an infinite number of follower types and extends prior work on learning in Stackelberg games without contextual information [26].
4.2 Learning with Contexts
We now shift to analyzing the setting where the leader has access to a data set consisting of tuples of contexts and follower types, $S_{m}=\{(\mathbf{z}_{1},f_{1}),\ldots,(\mathbf{z}_{m},f_{m})\}$. Given new contextual information, the leader aims to choose a strategy that maximizes her payoff against a follower, whose type is unknown. Here we introduce the study of the statistical complexity of distributional learning in Stackelberg games with side information.
Consider a distribution $\mathscr{D}$ over $\pazocal{Z}\times\mathscr{K}$. We define the leader’s payoff as a function of a policy $\pi:\pazocal{Z}\to\Delta(\pazocal{A})$ with respect to $\mathscr{D}$ as:
$u_{\mathscr{D}}(\pi)=\mathop{\mathbb{E}}_{(\mathbf{z},f^{(i)})\sim\mathscr{D}}\left[u(\mathbf{z},\pi(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi(\mathbf{z})))\right]$
and the leader’s average payoff on a sample $S_{m}=\{(\mathbf{z}_{1},f_{1}),\ldots,(\mathbf{z}_{m},f_{m})\}\in(\pazocal{Z}\times\mathscr{K})^{m}$ as:
$u_{S_{m}}(\pi)=\frac{1}{m}\sum_{i=1}^{m}\left[u(\mathbf{z}_{i},\pi(\mathbf{z}_{i}),b_{f_{i}}(\mathbf{z}_{i},\pi(\mathbf{z}_{i})))\right].$
The goal of the learner is to deploy an algorithm, $\textsc{Alg}:\cup_{m=0}^{\infty}(\pazocal{Z}\times\mathscr{K})^{m}\to\Delta(A)^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, the algorithm outputs a policy $\pi_{\textsc{Alg}}:\pazocal{Z}\to\Delta(A)$ that maximizes $u_{\mathscr{D}}(\pi)$.
Suppose that the leader has access to a set of hypotheses, $\pazocal{H}$, such that each hypothesis in $\pazocal{H}$, maps contexts to follower types ($h:\pazocal{Z}\to\mathscr{K}$). We define the true error function of a hypothesis $h$ with respect to a distribution, $\mathscr{D}$, as:
$\text{Err}_{\mathscr{D}}(h)=\mathbb{P}_{(\mathbf{z},f)\sim\mathscr{D}}\left[h(\mathbf{z})\neq f\right].$
Similarly, we define the empirical error of a hypothesis $h$ on a sample $S_{m}$ as:
$\text{Err}(h)=\frac{1}{m}|\{i\in[m]:h(\mathbf{z}_{i})\neq f_{i}\}|$
and the approximation error of a hypothesis class $\pazocal{H}$ on a distribution $\mathscr{D}$ to be:
$\text{Err}_{\mathscr{D}}(\pazocal{H}):=\inf_{h\in\pazocal{H}}\text{Err}_{\mathscr{D}}(h).$
Consider an optimal hypothesis, $h^{*}\in\pazocal{H}$, such that given any context, $h^{*}$ always predicts the follower type correctly, i.e. $\text{Err}_{\mathscr{D}}(h^{*})=0$. Since we are in the realizable setting, for the remainder of this section we will overload notation and use $\mathscr{D}$ to refer to the marginal distribution over the context space $\pazocal{Z}$.
We now define the Natarajan dimension, a measure of complexity of a multiclass hypothesis class (introduced by  Natarajan [29]), which will be useful in our setting. Informally, the Natarajan dimension measures the ability of a class of functions, whose image is a finite set, to fit complex patterns. [^4] Note that for binary labels, the definition of Natarajan dimension recovers the well-studied VC-dimension of a hypothesis class.
Definition 4.6 (Natarajan Dimension).
Let $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$ be a hypothesis class and let $S\subseteq\pazocal{Z}$. We say that $\pazocal{H}$ N-shatters $S$ if there exist $\phi_{1},\phi_{2}:S\to\mathscr{K}$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that:
$\forall\mathbf{z}\in T,\;g(\mathbf{z})=\phi_{1}(\mathbf{z})\quad\text{and}\quad\forall\mathbf{z}\in S\setminus T,\;g(\mathbf{z})=\phi_{2}(\mathbf{z}).$
The Natarajan dimension of $\pazocal{H}$, denoted $d_{N}(\pazocal{H})$, is the maximal cardinality of a set that is shattered by $\pazocal{H}$.
We define the realizable sample complexity of a Stackelberg-learning algorithm Alg for a hypothesis class $\pazocal{H}$ to be the function $m_{\textsc{Alg},\pazocal{H}}^{r}$ such that for every $\epsilon,\delta>0$, $m_{\textsc{Alg},\pazocal{H}}^{r}$ is the minimal integer such that for every $m\geq m_{\textsc{Alg},\pazocal{H}}^{r}$ and every distribution $\mathscr{D}$ over $\pazocal{Z}$, we have:
$\mathbb{P}_{S_{m}}\left(u_{\mathscr{D}}(\textsc{Alg}(S_{m}))<u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon\right)\leq\delta.$ (2)
where $\pi_{h^{*}}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$.
4.2.1 Vanilla ERM learner
We define an Empirical Risk Minimization (ERM) learner for the hypothesis class $\pazocal{H}$ as an algorithm, such that for any sample $S_{m}$ returns a function that minimizes the empirical error of that function relative to any other function in $\pazocal{H}$. Formally, we say:
$h_{ERM}$ $\in\arg\min_{h\in\pazocal{H}}\frac{1}{m}\sum_{i=1}^{m}\mathbbm{1}\{h(\mathbf{z}_{i})\neq f_{i}\}.$ (3)
We now show that the finiteness of the Natarajan dimension is a sufficient condition for learning the optimal policy for the leader in structured Stackelberg games. This can be achieved by using a simple ERM learner as in (3).
Proposition 4.7.
Consider an algorithm $\textsc{Alg}_{\text{ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times\mathscr{K})^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$ that outputs the policy:
$\pi_{\text{ERM}}(\mathbf{z})$ $\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h_{ERM}(\mathbf{z})}(\mathbf{z},\mathbf{x})),$
where $h_{ERM}$ satisfies equation (3). Then for any such $h_{\text{ERM}}$, there exists constant $C>0$, such that for every $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$,
$m^{r}_{\pazocal{A}_{\text{ERM}}}\leq C\left(\frac{d_{N}(\pazocal{H})\ln(K)\ln(\frac{1}{\epsilon})+\ln(\frac{1}{\delta})}{\epsilon}\right)$
We present the proof of Proposition 4.7 in Appendix B.
4.2.2 Stackelberg ERM learner
In this section, we show how to improve upon the sample complexity bound shown in Proposition 4.7, by leveraging the interaction between the Stackelberg game and the hypothesis class. Consider the following hypothesis class of possible leader policies:
$\Pi_{\pazocal{H},\pazocal{G}}=\Bigl{\{}\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})|h\in\pazocal{H},\pi_{h}(\mathbf{z})=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))\Bigl{\}}$
and the following variation of an ERM rule.
Definition 4.8 (Stackelberg ERM).
Consider a hypothesis class $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$ and a Stackelberg game, $\pazocal{G}$. We define the Stackelberg ERM learning algorithm, $\textsc{Alg}_{\text{S-ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times\mathscr{K})^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, outputs the policy $\pi_{\text{S-ERM}}:\pazocal{Z}^{m}\to\Delta(\pazocal{A})$ that satisfies:
$\pi_{\text{S-ERM}}\in\arg\min_{\pi\in\Pi_{\pazocal{H},\pazocal{G}}}\frac{1}{m}\sum_{i=1}^{m}\mathbbm{1}\{\pi(\mathbf{z}_{i})\neq\mathbf{x}_{i}^{*}\},$ (4)
where $\mathbf{x}_{i}^{*}=\arg\max_{x\in\Delta(\pazocal{A})}\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{i},\mathbf{x},b_{f_{i}}(\mathbf{z}_{i},\mathbf{x}))$.
We define the following measure of complexity of a hypothesis class with respect to a Stackelberg game.
Definition 4.9 (Stackelberg-Natarajan Dimension).
Consider a hypothesis class $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$, a Stackelberg game, $\pazocal{G}$, and let $S\subseteq\pazocal{Z}$. We say that $\pazocal{H}$ SN-shatters $S$ with respect to game $\pazocal{G}$, if there exist $\phi_{1},\phi_{2}:S\to\Delta(\pazocal{A})$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that:
$\forall\mathbf{z}\in T,\;\pi_{g}(\mathbf{z})=\phi_{1}(\mathbf{z})\quad\text{and}\quad\forall\mathbf{z}\in S\setminus T,\;\pi_{g}(\mathbf{z})=\phi_{2}(\mathbf{z}).$
The Stackelberg-Natarajan dimension of $\pazocal{H}$ with respect to game $\pazocal{G}$, denoted as $d_{SN,\pazocal{G}}(\pazocal{H})$, is the maximal cardinality of a set that is SN-shattered by $\pazocal{H}$.
Theorem 4.10.
For any hypothesis class $\pazocal{H}$, $d_{SN,\pazocal{G}}(\pazocal{H})\leq d_{N}(\pazocal{H})$. Additionally, there exists a hypothesis class and a Stackelberg game with distinct follower types for which the strict inequality holds.
Proof.
(Claim 1.) Consider a hypothesis class, $\pazocal{H}$. It suffices to show that for any set $S\subseteq\pazocal{Z}$ SN-shattered by $\pazocal{H}$, $S$ is also N-shattered by $\pazocal{H}$. This holds because for any pair of points $\mathbf{z}_{1},\mathbf{z}_{2}\in S$, for which there are two different leader strategies consistent with some $g\in\pazocal{H}$, (and assuming a fixed tie-breaking rule) then it must be that $g(\mathbf{z}_{1})\neq g(\mathbf{z}_{2})$, i.e. that function predicts two different follower types for these contexts.
(Claim 2.) Consider the set $\pazocal{Z}=\{\mathbf{z}_{1},\mathbf{z}_{2}\}$ and the set of follower types $\{f^{(0)},f^{(1)},f^{(2)}\}$. Also consider the following hypothesis class $\pazocal{H}=\{h_{i}\}_{i\in[4]}$, where $h_{1}(\mathbf{z}_{1})=h_{4}(\mathbf{z}_{1})=f^{(0)}$, $h_{2}(\mathbf{z}_{1})=h_{3}(\mathbf{z}_{1})=f^{(1)}$, $h_{1}(\mathbf{z}_{2})=h_{3}(\mathbf{z}_{2})=f^{(2)}$, and $h_{2}(\mathbf{z}_{2})=h_{4}(\mathbf{z}_{2})=f^{(1)}$.
It is easy to verify that for $\phi_{1},\phi_{2}$, where $\phi_{1}(\mathbf{z}_{1})=f^{(0)},\;\phi_{1}(\mathbf{z}_{2})=f^{(1)},\;\phi_{2}(\mathbf{z}_{1})=f^{(1)}$, and $\phi_{2}(\mathbf{z}_{2})=f^{(2)}$, $\pazocal{H}$ shatters the set $\{\mathbf{z}_{1},\mathbf{z}_{2}\}$, therefore, $d_{N}(\pazocal{H})=2$.
Now consider a Stackelberg game, where for all contexts $\mathbf{z}$, $\pi_{*}^{(0)}(\mathbf{z})\neq\pi_{*}^{(2)}(\mathbf{z})$, while $\pi_{*}^{(1)}(\mathbf{z}_{1})=\pi_{*}^{(0)}(\mathbf{z}_{1})$, and $\pi_{*}^{(1)}(\mathbf{z}_{2})=\pi_{*}^{(2)}(\mathbf{z}_{2})$, where $\pi_{*}^{(i)}:\pazocal{Z}\to\Delta(\pazocal{A})$, denotes the optimal policy of the leader against a follower type $f^{(i)}$. Then notice that we always have $\pi_{h^{*}}(\mathbf{z}_{1})=\pi_{*}^{(0)}(\mathbf{z}_{1})$ and $\pi_{h^{*}}(\mathbf{z}_{2})=\pi_{*}^{(2)}(\mathbf{z}_{2})$. Therefore, $d_{SN,\pazocal{G}}(\pazocal{H})=0$. ∎
As in the online setting, depending on the hypothesis class and the Stackelberg game at hand, the gap between the Stackelberg-Natarajan dimension and the Natarajan dimension can be arbitrarily large. We illustrate this via the following example.
Example 4.11.
Consider the hypothesis class of conjunctions, where each hypothesis is defined over a set of Boolean variables, $\pazocal{Z}=\{z_{1},\cdots,z_{n}\}$, such that $\pazocal{H}_{\text{conj}}=\{h:\{0,1\}^{n}\to\{0,1\}|h(z)=\land_{i\in S}z_{i}\land_{j\in T}z_{j},S,T\subseteq[n],S\cap T=\emptyset\}$. Since each hypothesis outputs in binary, this hypothesis class could naturally be considered in a Stackelberg game with two follower types, $f^{(0)},f^{(1)}$. The Natarajan dimension in this case reduces to the well-studied notion of the VC-dimension, which is known to be equal to $n$ for $\pazocal{H}_{\text{conj}}$. At the same time, the $d_{SN,\pazocal{G}}(\pazocal{H}_{\text{conj}})$ will be upper bounded by the size of the set $\{z\in\{0,1\}^{n}|\pi_{*}^{(0)}(z)\neq\pi_{*}^{(1)}(z)\}$, which could be as small as $1$ for two distinct follower types.
Theorem 4.12.
Consider the Stackelberg ERM policy, $\pi_{\text{S-ERM}}$, as in Definition 4.9 and let $h_{\text{S-ERM}}$ be the corresponding hypothesis in $\pazocal{H}$. Then universally for any such policy, there exists constant $C$, such that for every $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$ and distribution $\mathscr{D}$,
$m^{r}_{\pazocal{A}_{\text{ERM}}}\leq C\frac{1}{\epsilon}\left(d_{SN}(\pazocal{H})\log\left(\frac{d_{SN}(\pazocal{H})\cdot K}{\epsilon}\right)+\log\left(\frac{1}{\delta}\right)\right)$
Proof Sketch: The proof follows from a similar argument as that of the sample complexity bounds of multiclass learning [17]. The only part of the proof that does not apply here is Natarajan’s lemma. We prove an appropriate generalization in Lemma 4.13. The full proof of Theorem 4.12 is presented in Appendix B.
Lemma 4.13.
$|\Pi_{\pazocal{H},\pazocal{G}}|\leq|\pazocal{Z}|^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$.
We present the complete proof of Lemma 4.13 in Appendix B.
Remark 4.14.
It is worth noting that even when there is only one context (i.e. $\pazocal{Z}$ is a singleton), our results in this section are incomparable with those in Section 4.1. This is because in this section, the goal of the leader is to learn a deterministic mapping from contexts to follower types, assuming that one exists. So, if the contexts are the same between all examples in the training set, the follower types corresponding to these examples will also be the same. In contrast, the context-free setting we consider in Section 4.1 allows for a distribution over follower types. Thus the two bounds are generally incomparable.
5 Conclusion
In this work we study structured Stackelberg games, where contextual information is predictive of the follower’s type. The follower type that the leader is playing against is unknown, while the context that carries information about the state of the world is presented to both players. We show that the leader can use the unknown, fixed mapping from contexts to follower types to inform her strategy. In the online setting, our algorithms learn the optimal policy for the leader and achieve tight no-regret bounds. These bounds scale with a new notion of statistical complexity that directly exploits the structure of Stackelberg games. We show similar results in the distributional setting, where contexts and follower types are drawn from a (joint) distribution.
An interesting direction for future research is to analyze structured Stackelberg games in the beyond worst-case analysis framework. It has recently been shown that smooth online learning is as easy as distributional learning, with regret bounds scaling with the VC dimension of the hypothesis class. Extending these findings to multiclass learning and structured Stackelberg games might lead to improved results both from a statistical and computational point of view.
Appendix A Details and Proofs from Section 3
A.1 A technical subtlety
An important component of our analysis is that the leader must be able to compute the optimal strategy for a given context and follower type. For this, we state that for any context $\mathbf{z}\in\pazocal{Z}$ the leader incurs negligible regret by restricting herself to a subset $\mathscr{E}_{\mathbf{z}}\subseteq\Delta(\pazocal{A})$ which is finite and computable. To formally state our result we present the notion of a contextual best response region, which was first introduced in [24] and which we repeat here for completeness.
Definition A.1 (Contextual Follower Best-Response Region).
For context $\mathbf{z}$, follower type $f^{(i)}$, and follower action $a_{f}\in\pazocal{A}_{f}$, let $X_{\mathbf{z}}(f^{(i)},a_{f})\subseteq\pazocal{X}$ denote the leader’s mixed strategies such that a follower of type $f^{(i)}$ best-responds to all $x\in X_{\mathbf{z}}(f^{(i)},a_{f})$ by playing action $a_{f}$, that is, $X_{\mathbf{z}}(f^{(i)},a_{f})=\{\mathbf{x}\in\pazocal{X}:b_{f^{(i)}}(\mathbf{z},\mathbf{x})=a_{f}\}$
Notice that any possible follower best-response region is (1) a subset of $\Delta(\pazocal{A})$ and (2) the intersection of finitely-many halfspaces. Therefore, every $X_{\mathbf{z}}(f^{(i)},a_{f})$ is convex and bounded. However, the leader may not be able to optimize over the extreme points of a best-response region, as it may not be closed due to the followers’ tie-breaking.
Definition A.2 ($\delta$-approximate extreme points).
Fix a context $\mathbf{z}\in\pazocal{Z}$ and consider the set of all non-empty contextual best-response regions. For $\delta>0$, $\mathscr{E}_{\mathbf{z}}(\delta)$ is the set of leader mixed strategies such that for all best-response functions $\sigma$ and any $\mathbf{x}\in\Delta(\pazocal{A})$ that is an extreme point of cl$(\pazocal{X}_{\mathbf{z}}(\sigma))$, $\mathbf{x}\in\mathscr{E}_{\mathbf{z}}(\delta)$ if $\mathbf{x}\in\pazocal{X}_{\mathbf{z}}(\sigma)$. Otherwise there is some $\mathbf{x}^{\prime}\in\mathscr{E}_{\mathbf{z}}(\delta)$ such that $\mathbf{x}^{\prime}\in\pazocal{X}_{\mathbf{z}}(\sigma)$ and $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{1}\leq\delta$.
Notice that for sufficiently small $\delta$ i.e. $\delta=\mathscr{O}(\frac{1}{T})$, we get that the additional regret of the leader by choosing actions over $\mathscr{E}^{(i)}$ rather than $X(f^{(i)},a_{f})$ is negligible. Therefore, we use the shorthand $\mathscr{E}^{(i)}:=\mathscr{E}^{(i)}(\delta)$ for the rest of the section. The following lemma from [24] formalizes this notion.
Lemma A.3.
For any sequence of followers $f_{1},\ldots f_{T}$ and any leader policy $\pi$, there exists a policy $\pi^{(\mathscr{E})}:\pazocal{Z}\rightarrow\cup_{\mathbf{z}\in\pazocal{Z}}\mathscr{E}_{\mathbf{z}}$ that, when given context $\mathbf{z}$, plays a mixed strategy in $\mathscr{E}_{\mathbf{z}}$ and guarantees that $\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\pi^{(\mathscr{E})}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{(\mathscr{E})}(\mathbf{z}_{t})))\leq 1$. Moreover, the same result holds in expectation over any distribution over follower types.
Given Lemma A.3 and for ease of notation, in this paper we assume that the leader will always set the parameter $\delta$, such that it does not affect our regret bounds.
A.2 Using multiclass SOA as a black-box algorithm
Consider the following algorithm.
Algorithm 2 Learning with full-feedback and realizable follower types
1:for $t\in[T]$ do
2:     Observe $\mathbf{z}_{t}$
3:     Evaluate $\hat{f}_{t}=h_{t}(\mathbf{z}_{t})\leftarrow\text{SOA}(\mathbf{z}_{t})$
4:     Commit to $\mathbf{x}_{t}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{\hat{f}_{t}}(\mathbf{z}_{t},\mathbf{x}))$
5:     Observe follower type $f_{t}$
6:end
Here SOA refers to the online multiclass learning algorithm proposed by [17].
A.3 Online Learning in the Agnostic Setting
A.3.1 Upper bound
Algorithmic procedure.
Consider the “learning with expert advice” setting where the leader at each timestep $t$ would get to observe not only the context $\mathbf{z}_{t}$ but also $N$ advices, $\{\phi_{t}^{(1)},\cdots,\phi_{t}^{(N)}\}\in\Delta(\pazocal{A})^{N}$. Then we can denote the cumulative loss of expert $i$ at timestep $T$ as:
$L_{T,i}=-\sum_{t=1}^{T}u(\mathbf{z}_{t},\phi_{t}^{(i)},b_{f_{t}}(\mathbf{z}_{t},\phi_{t}^{(i)})).$
For an appropriately defined set of experts we can formalize the following algorithm.
Algorithm 3 Learning in the full-feedback, agnostic setting
1:input: time horizon $T$
2:Set $\eta=\sqrt{8\ln(N)/T}$
3:for $t\in[T]$ do
4:     Observe $\mathbf{z}_{t}$
5:     Receive advices $\{\phi_{t}^{(1)},\cdots,\phi_{t}^{(N)}\}\in\Delta(A)^{N}$
6:     Commit to $\mathbf{x}_{t}=\phi_{t}^{(i)}$ with probability proportional to $\exp(-\eta L_{t-1,i})$
7:     Receive utility $u(\mathbf{z}_{t},a_{l},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, where $a_{l}$ is drawn from $\mathbf{x}_{t}$
8:     Observe follower type $f_{t}$
9:end
See 3.15
Proof.
First consider the case where we construct an expert for every $h\in\mathscr{H}$ and keep track of their cumulative loss $L_{t,h}$ at each timestep $t$. In our case, for each $h\in\pazocal{H}$, we can construct an expert $i$ that will predict the type, $f_{t}^{(i)}=h(\mathbf{z}_{t})$ at that timestep and based on that evaluate:
$\phi_{t}^{(i)}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z}_{t},a_{l},b_{f_{t}^{(i)}}(\mathbf{z}_{t},\mathbf{x})).$
From the regret analysis of the Hedge algorithm (see [14] Chapter 2), we know that:
$\sum_{t\in[T]}\mathbb{E}\left[u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))\right]\geq\max_{i\in[N]}\sum_{t=1}^{T}u(\mathbf{z}_{t},\phi_{t}^{(i)},b_{f_{t}}(\mathbf{z}_{t},\phi_{t}^{i}))-\sqrt{\frac{1}{2}\ln(|\mathscr{H}|)T},$
where the expectation is taken over the randomness in the Hedge algorithm.
Now it remains to construct a smaller set of experts such that the upper bound on the regret depends on the $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, instead of $\ln(|\pazocal{H}|)$. For some time horizon $T$, we define $S_{T}=\{S\subset[T]:|S|\leq\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\}$.
For each $S\in S_{T}$ and $g:S\rightarrow\Delta(\pazocal{A})$, we construct an expert, $E_{S,g}$. Expert $E_{S,g}$ immitates an instance of the SSOA algorithm, where SSOA errs at all timesteps $t\in S$ and for each timestep $t\in S$, the optimal strategy for the leader at the timestep is given by $g(t)$. We show the implementation of each expert in Algorithm 4 below.
Algorithm 4 Implementation of expert $E_{S,g}$
1:input: time horizon $T$
2:Set $V_{0}=\pazocal{H}$
3:for $t\in[T]$ do
4:     Observe $\mathbf{z}_{t}$
5:     Let $\mathbf{x}^{(h,t)}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))\quad\forall h\in\pazocal{H}$
6:     Let $X_{t}=\{\mathbf{x}^{(h)}:h\in V_{t-1}\}$triangleright $|X_{t}|\leq K$
7:     Let $V_{t}^{(\mathbf{x})}=\{h\in V_{t-1}:\mathbf{x}^{(h,t)}=\mathbf{x}\}\quad\forall\mathbf{x}\in X_{t}$
8:     Play action $l_{t}=\arg\max_{\mathbf{x}\in X_{t}}\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$
9:     Set $l_{t}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{\mathbf{z}_{t}\rightarrow k})$
10:     if $t\in S$ then
11:         Predict $g(t)$
12:         $V_{t+1}\leftarrow V_{t}^{\mathbf{z}_{t}\rightarrow g(t)}$
13:     else
14:         Predict $l_{t}$
15:         $V_{t+1}\leftarrow V_{t}^{\mathbf{z}_{t}\rightarrow l_{t}}$
16:end
Notice that for each possible timestep $t\in T$, a single context $\mathbf{z}_{t}$ is presented, so there are at most $K$ possible optimal strategies for the leader, one for each follower type. So, the number of experts that can be constructed in this way is:
$\sum_{i=0}^{\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})}\binom{T}{i}(K-1)^{i}\leq(TK)^{\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})}.$
So using the above regret upper bound with this set of experts we get that Algorithm 3 achieves utility at least:
$\max_{S,g}-L_{S,g,T}-\sqrt{\frac{1}{2}\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})T\ln(TK)},$
where $-L_{S,g,T}$ is the accumulated utility of expert $E_{S,g}$ up to timestep $T$. It remains to show that $\max_{S,g}-L_{S,g,T}\geq\max_{h\in\mathscr{H}}-L_{h,T}$. Notice that for any $h\in\mathscr{H}$ and for any set of timesteps $S$ that SSOA errs in, there exists an expert in the constructed set of experts such that $L_{S,g,T}=L_{h,T}$ when executed on the sequence $(\mathbf{z}_{1},h(\mathbf{z}_{1})),\cdots,(\mathbf{z}_{T},h(\mathbf{z}_{T}))$.
A.3.2 Lower bound
See 3.16
Proof.
We let $d=\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$. For simplicity assume that $T$ is an integer multiple of $d$, say $T=kd$ for $k\in\mathbb{N}$. Consider a tree, $(\mathbf{v}_{1},\mathbf{v}_{1},\ldots,\mathbf{v}_{2^{d}-1})$, of depth, $d$, that is SL-shattered by $\pazocal{H}$ such that each node $\mathbf{v}_{i}\in\pazocal{Z}$ has two edges (each leading to a child) corresponding to two strategies for the leader $\mathbf{x}_{i,1},\mathbf{x}_{i,2}\in\Delta(\pazocal{A})$, such that $\mathbf{x}_{i,1}\neq\mathbf{x}_{i,2}$. By definition, we have that for each $i$, there exist at least two functions $h_{i,1},h_{i,2}\in\pazocal{H}$ such that at context $\mathbf{v}_{i}$, they predict follower types $f_{1}=h_{i,1}(\mathbf{v}_{i})$ and $f_{2}=h_{i,2}(\mathbf{v}_{i})$, against which the strategies $\mathbf{x}_{i,1}$ and respectively $\mathbf{x}_{i,2}$ are optimal. In other words, we construct a Stackelberg game, playing the optimal strategy against the mispredicted follower type gives utility 00, whereas playing the optimal strategy against the correct follower type gives utility $1$ (for an explicit construction see proof of Theorem 3.8). Then, we have that for every $i$:
$u(\mathbf{v}_{i},\mathbf{x}_{i,1},b_{f_{1}}(\mathbf{v}_{i},\mathbf{x}_{i,1}))-u(\mathbf{v}_{i},\mathbf{x}_{i,2},b_{f_{1}}(\mathbf{v}_{i},\mathbf{x}_{i,2}))=1$
and similarly,
$u(\mathbf{v}_{i},\mathbf{x}_{i,2},b_{f_{2}}(\mathbf{v}_{i},\mathbf{x}_{i,2}))-u(\mathbf{v}_{i},\mathbf{x}_{i,1},b_{f_{2}}(\mathbf{v}_{i},\mathbf{x}_{i,1}))=1.$
We will now construct a sequence $(\mathbf{z}_{1},\mathbf{x}_{1}),(\mathbf{z}_{2},\mathbf{x}_{2}),\ldots,(\mathbf{z}_{T},\mathbf{x}_{T})$, which we will show forces any algorithm to meet the lower bound in the theorem statement. From the SL-shattered tree, we follow a root-to-leaf path, say $(\tilde{\mathbf{z}}_{1},\tilde{\mathbf{x}}_{1}),\ldots,(\tilde{\mathbf{z}}_{d},\tilde{\mathbf{x}}_{d})$. For each node $\tilde{\mathbf{z}}_{i}$ on the path, where $i\in[d]$, we construct a set of $k$ pairs of contexts and follower types $\{(\mathbf{z}_{(i-1)k+1}^{a},\mathbf{x}_{(i-1)k+1}^{a}),\ldots,(\mathbf{z}_{ik}^{a},\mathbf{x}_{ik}^{a})\}$ as follows.
  • We set $\mathbf{z}_{(i-1)k+1}^{a}=\ldots=\mathbf{z}_{ik}^{a}=\tilde{\mathbf{z}}_{i}$.
  • We draw each $\mathbf{x}_{t}^{a}$ for $t\in\{(i-1)k+1,\ldots,ik\}$, independently and uniformly at random from the two optimal strategies corresponding to the edges leading to the two children of $\mathbf{z}_{i}$. We denote the follower type and optimal strategy for each child, $f_{A,i},\mathbf{x}_{A,i}$ and $f_{B,i},\mathbf{x}_{B,i}$ respectively.
For each set of $k$ examples, let $T_{i}=\{(i-1)k+1,\ldots,ik\}$ be the set of indeces of the examples corresponding to $\tilde{\mathbf{z}}_{i}$. It suffices to lower bound:
$\max_{h\in\pazocal{H}}\sum_{t\in T_{i}}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$
For ease of notation, we denote $h^{*}=\arg\max_{h\in\pazocal{H}}\sum_{t\in T_{i}}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))$, $\;u_{h^{*},t}=u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))$, $u_{t}=u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, and $u_{t}^{*}=u(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t})))$, where $\pi_{f_{t}}(\mathbf{z}_{t})$ corresponds to the optimal policy against the true follower type and context at time $t$. So equivalently, we want to lower bound:
$\sum_{t\in T_{i}}u_{h^{*},t}-u_{t}=\sum_{t\in T_{i}}(u_{t}^{*}-u_{t})-(u_{t}^{*}-u_{h^{*},t}).$
Let $r_{i}$ be the number of timesteps in $T_{i}$, where $f_{A,i}$ appears in the example. Then:
$u_{t}^{*}-u_{h^{*},t}=\begin{cases}k-r\quad\text{if}\;r\geq k/2\\ r\hskip 28.45274pt\text{if}\;r<k/2\end{cases}$
Therefore, we get that $k/2-\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}=|r-k/2|$. By taking an expectation over the randomness in the $f_{t}$’s, we get that:
$k/2-\mathbb{E}\left[\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}\right]=\mathbb{E}\left[|r-k/2|\right]\geq\sqrt{k/8},$
where the last inequality holds by Khinchine’s inequality (see Lemma A.9 in [14]). Now it suffices to sum over all $i\in[d]$ to get:
$\frac{dk}{2}-\mathbb{E}\left[\sum_{t\in[T]}u_{t}^{*}-u_{h^{*},t}\right]=\sum_{i\in[d]}\left(\frac{k}{2}-\mathbb{E}\left[\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}\right]\right)\geq d\sqrt{k/8}$
Finally, notice that $dk/2=T/2=\sum_{t\in[T]}(u_{t}^{*}-u_{t})$ for any algorithm, which concludes the proof. ∎
Appendix B Details and Proofs from Section 4
See 4.7
Proof.
We have that:
$\mathbb{P}_{S_{m}}\left[u_{\mathscr{D}}(\pi_{h^{*}})-u_{\mathscr{D}}(\pi_{\text{ERM}})\geq\epsilon\right]$
$\leq$ $\mathbb{P}_{S_{m}}\left[\mathop{\mathbb{E}}_{\mathbf{z}\sim\mathscr{D}}\left[u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z})))\right]>\epsilon\right]$
By the boundedness of the utility function we have that the expectation above lies within $[-1,1]$. Therefore, we get:
$\leq$ $\mathbb{P}_{S_{m}}\left[\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z})))>0\right]>\epsilon\right]$
But getting less utility than the optimal predictor in the class, implies that our prediction of the follower’s type is incorrect. So,
$\leq$ $\mathbb{P}_{S_{m}}\left[\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[h_{*}(\mathbf{z})\neq h_{\text{ERM}}(\mathbf{z})\right]>\epsilon\right]$
But from Theorem 6 in [17], we know that for any $\epsilon,\delta>0$, there exists constant $C_{2}$ such that for any $m\geq C_{2}\left(\frac{d_{N}(\pazocal{H}))\ln(K)\ln(\frac{1}{\epsilon})+\ln(\frac{1}{\delta})}{\epsilon}\right)$, we have that for any $h_{\text{ERM}}$ that satisfies equation (3):
$\mathbb{P}_{S_{m}}\left(\text{Err}_{\mathscr{D}}(h_{\text{ERM}})>\text{Err}_{\mathscr{D}}(\pazocal{H})+\epsilon\right)\leq\delta$
Since we are in the realizable setting, we know that $\text{Err}_{\mathscr{D}}(\pazocal{H})=0$. By definition, we have that $\text{Err}_{\mathscr{D}}(h_{\text{ERM}})>\epsilon\iff\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[h_{\text{ERM}}(\mathbf{z})\neq h^{*}(\mathbf{z})\right]>\epsilon$, which concludes our proof. ∎
B.1 Proof of Theorem 4.12
See 4.12
Proof.
Note that the following proof follows parts of Balcan [3], which we reproduce for completeness. By definition, it suffices to show that for any $\epsilon,\delta\in(0,1)$ if we draw a sample $S$ from $\mathscr{D}$ of size:
$m>C\cdot\frac{1}{\epsilon}\Bigr{[}d_{SN,\pazocal{G}}(\pazocal{H})\log_{2}\Bigl{(}\frac{d_{SN,\pazocal{G}}(\pazocal{H})\cdot K}{\epsilon}\Bigl{)}+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
for some large enough constant $C$, then with probability at least $1-\delta$,
$\mathbb{P}_{S_{m}}\left(u_{\mathscr{D}}(\pi_{\text{S-ERM}})\geq u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon\right)$ (5)
where $\pi_{h^{*}}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$. In other words, it suffices to prove that all the policies in $\Pi_{\pazocal{H},\pazocal{G}}$ with $u_{\mathscr{D}}(\pi_{h})<u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon$ have low utility in the training sample too, that is, $u_{S}(\pi_{h})\leq u_{S}(\pi_{h^{*}})$.
It suffices then to bound the following bad event:
$B:\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}\text{ with }u_{S}(\pi_{h})=u_{S}(\pi_{h^{*}})\text{ but }u_{D}(\pi_{h})<u_{D}(\pi_{h^{*}})-\epsilon.$
Let us denote the training sample by $S=\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{m}\}$. Now suppose $S^{\prime}=\{\mathbf{z}^{\prime}_{1},\mathbf{z}^{\prime}_{2},\ldots,\mathbf{z}^{\prime}_{m}\}$ is another sample drawn i.i.d. from $\mathscr{D}$ (a ghost sample). We define the event:
$B^{\prime}:\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}\text{ with }u_{S}(\pi_{h})=u_{S}(\pi_{h^{*}})\text{ but }u_{S^{\prime}}(\pi_{h})<u_{S^{\prime}}(\pi_{h^{*}})-\epsilon/2.$
Lemma B.1.
If $m>\frac{8}{\epsilon}$, then $\mathbb{P}[B]\leq 2\mathbb{P}[B^{\prime}]$.
Proof.
Suppose that we know that there exists an $\pi_{h}$ that is consistent with $S$ but $L_{\mathscr{D}}(\pi_{h}):=u_{D}(\pi_{h^{*}})-u_{D}(\pi_{h})>\epsilon$. Since $S^{\prime}$ is drawn i.i.d. from $\mathscr{D}$, $\mathbb{E}[L_{S^{\prime}}(\pi_{h})]\geq\epsilon m$. By a Chernoff bound, $\mathbb{P}[L_{S^{\prime}}(\pi_{h})\leq\epsilon m/2]<e^{-m\epsilon/8}\leq\frac{1}{2}$ for $m>\frac{8}{\epsilon}$. This implies that:
$\frac{\mathbb{P}[B^{\prime}]}{\mathbb{P}[B]}\geq\frac{\mathbb{P}[B^{\prime}\cap B]}{\mathbb{P}[B]}=\mathbb{P}[B^{\prime}|B]\geq\frac{1}{2},$
so $\mathbb{P}[B]\leq 2\mathbb{P}[B^{\prime}]$. ∎
Thus, it suffices to bound $\mathbb{P}[B^{\prime}]$ (this probability is over the choices of $S$ and $S^{\prime}$). Given two samples $S$ and $S^{\prime}$, consider the random process SwapR:
  • For $i$ from $1$ to $m$, flip a fair coin.
  • If heads, swap $\mathbf{z}_{i}$ and $\mathbf{z}^{\prime}_{i}$, else do nothing.
Let $T$ and $T^{\prime}$ denote the resulting collections.
Lemma B.2.
Suppose $S$ and $S^{\prime}$ are chosen according to $\mathscr{D}$ and SwapR is applied. Then $T$ and $T^{\prime}$ are identically distributed to $S$ and $S^{\prime}$.
Now, we define the event:
$B^{\prime\prime}:\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}\text{ with }L_{T}(\pi_{h})=0\text{ but }L_{T^{\prime}}(\pi_{h})>\epsilon/2.$
By Claim B.2, $\mathbb{P}[B^{\prime\prime}]=\mathbb{P}[B^{\prime}]$ so, it suffices to upper bound $\mathbb{P}[B^{\prime\prime}]$. Notice that the first probability is over the choice of $S$, $S^{\prime}$, and the randomness of SwapR, while the second is over the choice of $S$ and $S^{\prime}$.
Lemma B.3.
Fix $h\in\pazocal{H}$. Then $\mathbb{P}[L_{T}(\pi_{h})=0\land L_{T^{\prime}}(\pi_{h})>\epsilon/2|S,S^{\prime}]\leq 2^{-\epsilon m/2}.$
Proof.
Consider:
$\pi_{h}(\mathbf{z}_{1}),\pi_{h}(\mathbf{z}_{2}),\ldots,\pi_{h}(\mathbf{z}_{m})$
$\pi_{h}(\mathbf{z}^{\prime}_{1}),\pi_{h}(\mathbf{z}^{\prime}_{2}),\ldots,\pi_{h}(\mathbf{z}^{\prime}_{m}).$
Notice the following: (1.) If there is a column, with both predictions wrong, then $L_{T}(\pi_{h})=0$ can never happen, so we are done. (2.) Similarly, if more than $(1-\epsilon/2)m$ columns have both predictions correct, $L_{T^{\prime}}(\pi_{h})>\epsilon/2$ cannot happen.
So, there must be at least $r\geq\epsilon m/2$ columns with one correct and one wrong prediction. In this case, SwapR must always ensure that the correct prediction remains on top, which happens with probability at most $2^{-r}\leq 2^{-\epsilon m/2}$. ∎
We are now ready to upper bound $\mathbb{P}[B^{\prime\prime}]$.
Lemma B.4.
$\mathbb{P}[B^{\prime\prime}]\leq\Pi_{\pazocal{H},\pazocal{G}}[2m]\cdot 2^{-\epsilon m/2}$.
Proof.
By definition, we have that:
$\mathbb{P}[B^{\prime\prime}]$ $=\mathop{\mathbb{E}}_{S,S^{\prime}}[\mathbb{P}_{\texttt{SwapR}}[\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}},\;L_{T}(h)=0\land L_{T^{\prime}}(h)\geq m\epsilon/2|S,S^{\prime}]]$
$=\mathop{\mathbb{E}}_{S,S^{\prime}}[\mathbb{P}_{\texttt{SwapR}}[\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}[S\cup S^{\prime}],\;L_{T}(h)=0\land L_{T^{\prime}}(h)\geq m\epsilon/2|S,S^{\prime}]].$
Using the union bound, we get that:
$\mathbb{P}[B^{\prime\prime}]$ $\leq\mathop{\mathbb{E}}_{S,S^{\prime}}\Bigr{[}\sum_{\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}[S\cup S^{\prime}]}\mathbb{P}_{\texttt{SwapR}}[\;L_{T}(\pi_{h})=0\land L_{T^{\prime}}(\pi_{h})\geq m\epsilon/2|S,S^{\prime}]\Bigr{]}$
$\leq\Pi_{\pazocal{H},\pazocal{G}}[2m]2^{-\epsilon m/2}$
Combining all these, we get that $\mathbb{P}[B]\leq\delta$, whenever $2\Pi_{\pazocal{H},\pazocal{G}}[2m]2^{-\epsilon m/2}\leq\delta$, which proves that:
$m>\frac{2}{\epsilon}\Bigr{[}\log_{2}(2\Pi_{\pazocal{H},\pazocal{G}}[2m])+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
It remains to show that $\Pi_{\pazocal{H},\pazocal{G}}\leq d_{SN,\pazocal{G}}(\pazocal{H})\ln(K)$. We prove this using the Lemma 4.13, which states that $\Pi_{\pazocal{H},\pazocal{G}}[m]\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$. So we get that:
$m$ $>\frac{2}{\epsilon}\Bigr{[}\log_{2}(2\cdot(2m)^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})})+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
$=\frac{2}{\epsilon}\Bigr{[}1+d_{SN,\pazocal{G}}(\pazocal{H})\log_{2}(m)+d_{SN,\pazocal{G}}(\pazocal{H})+2d_{SN,\pazocal{G}}(\pazocal{H})\log(K)+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
Rearranging, we obtain the requirement that:
$m>\frac{2d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigr{[}\log_{2}(m)\Bigr{]}+\frac{2}{\epsilon}\Bigr{[}2d_{SN,\pazocal{G}}(\pazocal{H})\log(K)+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}+d_{SN,\pazocal{G}}(\pazocal{H})+1\Bigr{]}$
Using lemma B.5, a sufficient condition for the above to hold is that:
$m>\frac{8d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigr{[}\log_{2}\Bigl{(}\frac{4d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigl{)}\Bigr{]}+\frac{4}{\epsilon}\Bigr{[}2d_{SN,\pazocal{G}}(\pazocal{H})\log(K)+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}+d_{SN,\pazocal{G}}(\pazocal{H})+1\Bigr{]}$
And a sufficient condition for the above is:
$m>\frac{8d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigr{[}\log_{2}\Bigl{(}\frac{4d_{SN,\pazocal{G}}(\pazocal{H})\cdot K}{\epsilon}\Bigl{)}\Bigr{]}+\frac{4}{\epsilon}\Bigr{[}\log\Bigl{(}\frac{1}{\delta}\Bigl{)}+d_{SN,\pazocal{G}}(\pazocal{H})+1\Bigr{]}$
The above holds if there exists a constant $C$, large enough such that:
$m>C\cdot\frac{1}{\epsilon}\Bigr{[}d_{SN,\pazocal{G}}(\pazocal{H})\log_{2}\Bigl{(}\frac{d_{SN,\pazocal{G}}(\pazocal{H})\cdot K}{\epsilon}\Bigl{)}+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}.$
Lemma B.5 (Shalev-Shwartz and Ben-David [34], Lemma A.2).
Let $a\geq 1$ and $b>0$. Then $x\geq 4a\log(2a)+2b$ implies that $x\geq a\log(x)+b$.
B.2 Proof of Lemma 4.13
See 4.13
Proof.
We will show that for $|Z|=m$:
$|\Pi_{\pazocal{H},\pazocal{G}}|$ |$\leq|\{B\subseteq\pazocal{Z}:\pazocal{H}\;\text{SN-shatters}\;B\}|\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$ (6)
$\leq\sum_{i=1}^{d_{SN,\pazocal{G}}(\pazocal{H})}\binom{m}{i}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$ (7)
$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$ (8)
Notice that equation (7) holds by enumerating all possible subsets of $\pazocal{Z}$ that can be SN-shattered by $\pazocal{H}$ up to $d_{SN,\pazocal{G}}(\pazocal{H})$, which is the largest size subset that $\pazocal{H}$ can shatter. So it suffices to prove equation (6). We proceed by induction on the size of $|\pazocal{Z}|$.
Base Case: For $|\pazocal{Z}|=1$, we have that $|\Pi_{\pazocal{H},\pazocal{G}}|\leq K$ so (6) holds.
Inductive Hypothesis: For $|\pazocal{Z}|=m$ assume that (6) holds.
Inductive Step: We will show (6) for $|S|=m+1$. Consider the set $\pazocal{Z}=\{\mathbf{z}_{0},\mathbf{z}_{1},\cdots,\mathbf{z}_{m}\}$ and we define the following notation:
$\Lambda^{(i)}=\{\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}|\pi_{h}(\mathbf{z}_{0})=i\}$
$\Lambda^{(i,j)}=\{\pi_{h}\in\Lambda^{(i)}|\exists\pi_{h}^{\prime}\in\Lambda^{(j)}\;\text{s.t.}\;\pi_{h}(\mathbf{z})=\pi_{h}^{\prime}(\mathbf{z})\;\forall\mathbf{z}\in\pazocal{Z}\setminus\{\mathbf{z}_{0}\}\}$
For $i,j\in[K]$ also denote $\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}$ to be all the policies $\pi_{f}:\pazocal{Z}\setminus\{\mathbf{z}_{0}\}\to[K]$ that can be extended to a policy in $\Pi_{\pazocal{H},\pazocal{G}}$ both by defining $\pi_{f}(\mathbf{z}_{0})=i$ and by defining $\pi_{f}(\mathbf{z}_{0})=j$. Then we claim the following:
$|\Pi_{\pazocal{H},\pazocal{G}}|\leq|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}|\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|,$ (9)
where $|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|$ denotes the restriction of $\Pi_{\pazocal{H},\pazocal{G}}$ to the set $\pazocal{Z}\setminus\{\mathbf{z}_{0}\}$.
(Proof of equation (9)). Consider the $\ell$’th labeling of elements in $\pazocal{Z}\setminus\{\mathbf{z}_{0}\}$ that is consistent with a policy in $\Pi_{\pazocal{H},\pazocal{G}}$. Let $\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}$ denote the policies in $\Pi_{\pazocal{H},\pazocal{G}}$ that are consistent with that labeling. Then we have that:
$|\Pi_{\pazocal{H},\pazocal{G}}|$ $=\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}$
$=\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}1+\left(\sum_{i=1}^{K}\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}\right)-1$
$=|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}-1$
$\leq|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}-1+\sum_{j=1}^{K}\mathbbm{1}\{\exists\pi,\pi^{\prime}\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\wedge\pi^{\prime}(\mathbf{z}_{0})=j\}$
$=|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\left(\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}-1\right)$
$\hskip 85.00012pt+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}\mathbbm{1}\{\exists\pi,\pi^{\prime}\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\wedge\pi^{\prime}(\mathbf{z}_{0})=j\}$
$\leq|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}\mathbbm{1}\{\exists\pi,\pi^{\prime}\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\wedge\pi^{\prime}(\mathbf{z}_{0})=j\}$
$=|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}|\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|.$ (10)
We proceed with the induction. From the inductive hypothesis we have that:
$|\displaystyle|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|$$\leq|\{B\subseteq\pazocal{Z}\setminus\{\mathbf{z}_{0}\}:\pazocal{H}\;\text{SN-shatters}\;B\}|\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$
$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})},$ (11)
where the last inequality holds from equation (8). Also, notice that each $\Lambda^{(i,j)}$ SN-shatters $\mathbf{z}_{0}$ by definition. So $\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{x_{0}\}}$ cannot shatter a set larger than $d_{SN,\pazocal{G}}(\pazocal{H})-1$. Therefore, by the inductive hypothesis we get that:
$|\displaystyle|\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{x_{0}\}}|$$\leq|\{B\subseteq\pazocal{Z}\setminus\{\mathbf{z}_{0}\}:\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}\;\text{SN-shatters}\;B\}|\cdot K^{2(d_{SN,\pazocal{G}}(\pazocal{H})-1)}$
$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}\cdot K^{2(d_{SN,\pazocal{G}}(\pazocal{H})-1)}$ (12)
Combining inequalities (10), (11), and (12), we get that:
$|\Pi_{\pazocal{H},\pazocal{G}}|$ $\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}\cdot K^{2(d_{SN,\pazocal{G}}(\pazocal{H})-1)}$
$=m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}+m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$
$=m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}K^{2d_{SN,\pazocal{G}}(\pazocal{H})}(m+1)$
$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$